PIP-II report for April 20, 2018

Highlights

  • ANL successfully qualified HWR cavity #3.
  • Qualified SSR1 cavities S113 and S114 with low power RF coupler at STC.
  • PIP2IT continues to increase beam power (now at 3.5 ms, 5mA, 20 Hz); construction of the east wall of cave was completed, preparing for cryo transfer line piping; facility was visited by Steve Binkley on April 20th.
  • A/E Acquisition Plan Corrections Underway.

ESH&Q (Teri Dykhuis, John Anderson)

  • Concerning the Wetlands Non-Jurisdictional Determination: By April 30, USACE will submit recommendations to an internal 21 day review. Targeting June 1st to deliver an official response.
  • Working with Fermilab Ecologist and Arcadis (our EA subcontractor) Biologist, on how to assess potential impact of PIP-II project to the Rusty Patched Bumble Bee, an endangered species as of 2017.
  • Another issue is the desire to relocate the proposed temporary stockpiles in order to prevent impact to the restored prairie.

Project Management (Marc Kaducak)

  • Development of the RLS continued with good progress in Safety Systems, Beam Instrumentation, and Magnets/Power Supplies.
  • Based on feedback from the L2 managers, we are updating the estimates for resources required to prepare for CD-2/3a.
  • Work on reorganizing the WBS dictionary to map to the WBS continues. We will be presenting the concept to the L2 managers.

Technical Integration (Arkadiy Klebaner)

  • Finalizing development of the Systems Engineering Management Plan and Design Reviews Plan
  • Working with L2 managers to define CD2 Design Readiness Plan
  • Started development of Design Deliverables Document for CD2
  • Requirements Verification and Interface Control Documents work proceeding
  • Continue development of the outline for the Preliminary Design Report
  • Members of the instrumentation and accelerator physics groups met with representatives from the University of Michigan to discuss extinction measurements at PIP2IT.  The Mu2e-II proposal is looking for a well-defined bunch pattern with 10-11 extinction over a 1.6 microsecond interval compared to beam during a 100 nsec interval.  The Michigan group is interested in working towards measuring the relative intensity of beam in occupied buckets vs beam in unoccupied buckets, as defined by the MEBT chopper.

International Coordination (Steve Holmes, Lia Merminga)

  • Welding of SSR1 dressed cavity S104 initiated at BARC. On track for shipment by end of May.
  • RFPI TRS Rev 5 released in TeamCenter.
  • Job descriptions for next round of India/DAE short-term visitors under development.
  • Final shipment of warm magnets for PIP2IT is being packaged at BARC. Expect shipment week of April 23.
  • Planning workshops with CEA and IN2P3 in France at the end of June.

SRF and Cryo Systems (Allan Rowe, Genfa Wu)

HWR

  • ANL successfully qualified the HWR #3

  • New 7kW TOMCO solid-state amplifier with 10 kW circulator was delivered and installed at ANL.

  • HWR coupler multipactor studies completed at Fermilab. Recommendation: increase the DC-bias voltage up to + 1.5 kV to suppress multipactor on the cold window.

  • The work on updating HWR CM Interface Control Documents in the TeamCenter is ongoing.
  • FNAL agreed to keep the HWR BPMs rotated by 45 degrees with respect to X&Y planes.

SSR1

  • Qualified SSR1 cavity S113 and S114 with low power RF coupler at STC.
  • Cavity S111 received the vacuum-end high power RF coupler (CPI-05) at MP9
  • Cavity S113 is at MP9 to receive the vacuum-end high power RF coupler (CPI-01)
  • Qualified two CPI vacuum-end high power RF couplers (CPI-01, CPI-03) on the RF test stand, at room temperature, with 10kW power in full reflection mode.
  • Procurement of 6 vacuum-end couplers was approved. First two units are expected by mid-June.
  • Design and drawings of lifting tooling for string assembly are completed. Next: procurement
  • Process for procurement of thermal shield and cooldown warm up line for the coldmass was started.
  • Lab 2 cleanroom: working on the preparation for the qualification/certification.

LB650

  • New mechanical model of dressed cavity being analyzed by VECC to improve mechanical performance under room temperature loading conditions.

HB650

  • Preliminary magnetic shielding analysis presented to 650 MHz collaboration meeting.

  • Cryomodule design status and assembly strategy presented at 650 MHz collaboration meeting.
  • First photos of prototype 650 MHz high power RF couplers arrived from vendor.

Cryogenics Plant

  • Met with ESH&Q staff to discuss Cryoplant Building (CPB) ODH assessment need by dates
  • Scheduled meeting with Conventional Facilities (CF) Manager to discuss CPB room data sheets (RDS) and Technical Requirements Specification (TRS)
  • Reviewed and responded to BARC CP technical evaluation. Asked BARC to request vendors for additional information: 1. 20% increase in 2K capacity, upgrade path and associated cost proposal. 2. T-S diagram for 2K capacity of cryoplant with no shields flow.
  • Reviewed ICD requirements

Cryogenics Distribution System

  • Working on interface control document for the CDS.
  • Opened dialogue with ESS (European Spallation Source) colleagues to understand the ESS cryogenic distribution system design and requirements. The ESS CDS is very similar to the proposed PIP-II CDS design utilizing individual bayonet cans for interfacing to cryomodules.

Accelerator Systems (Paul Derwent)

High Power RF

  • Procurement of HWR amplifiers.
  • The 7kW, 325 MHz amplifier design is with the vendor, ECIL. A new version of the 40kW, 650 MHz amplifier is under review, before requesting approval by RRCAT management.
  • One of the two 75 kW SigmaPhi amplifiers for the RFQ continues to have internal communication errors, which affect our ability to control the amplifier.  At this time, the errors turn off the amplifier, so we are not in danger of damaging equipment.  We are in communication with the vendor about the problem but have not received information yet on how to address it.
  • MDB Test Infrastructure –
    • Coupler Test circulator 3-D model was received from vendor and incorporated into design library for 650 MHz parts.  A 3-D model for the circulator stand is in progress and other coax parts will be incorporated.  A 2-D drawing has been done to verify routing and parts.  We have identified the type of parts and approximate quantities to order for the coax transmission system.  Once the 3-D model is more complete we will have a meeting of all the stakeholders to confirm the design is on-track.  We will then soon follow with a formal review and subsequent placement of orders.
    • The STC portion is being designed, and will be reviewed soon.

Low level RF

  • LLRF components purchase requisitions for HWR, SSR1 and test stands.
  • Productive combined LLRF and BPM studies at PIP2IT.
  • IILFC LLRF chassis continued development.

Vacuum

  • The investigation of gas burst on RFQ (PIP2IT) concluded with more than 95% of bursts identified after month long RGA-Helium study. Concluded that 25% of them originate from Tuner plugs and 70% from pumping port blanks. A proposal of verification and correction is in progress.
  • Started working on the interface documents, as part of the CD2 requirements.

Linac Installation and Commissioning (Fernanda G. Garcia)

PIP2IT

  • PIP2IT continues to make progress in increasing the power running at PIP2IT: reached 3.5 ms at 5mA, 20 Hz. Investigating beam instabilities under these conditions.
  • Construction of the east wall at PIP2IT cave extension was completed. Continued with preparation on the cryo transfer line piping welding.
  • The facility was visited by Steve Binkley on April 20.  Also, during the week received visitors from University of Michigan and CERN.

Conventional Facilities (Steve Dixon)

  • A/E Acquisition Plan Corrections Underway – DOE has returned acquisition plan for the A/E re-compete with comments on Friday 20APR18.  A meeting with DOE is scheduled for Tuesday 24APR18 to go over the comments and present the strategy for A/E conventional facilities procurements.
  • VE Workshop Evaluations Returned –  Comments on the draft results from the A/E on the cost/schedule evaluations from the VE workshop were returned last week.  A phone meeting was held on 20APR18 to discuss the comments and related information.
  • P2MAC Recommendations Response entered in iTrack.
  • Site Prep Design – Two phone meetings were held with the A/E concerning the Site Prep design work.  Discussions included the requirements for the cost loading of the work into the RLS, quality control during design, milestones and information needed from Fermilab.  The kick-off meeting is being scheduled.